∫ (x)= ∫ sinx²y dx
1
Expert's answer
2020-04-21T14:39:11-0400

Let "I=\int sinx^2y dx"

Here y is constant ,Since integration is with respect to x .

"\implies I=\int sin (x\sqrt{y})^2."

Let "x\sqrt{y}=u" "\implies dx=\frac{1}{\sqrt{y}} du."

"\implies I=\frac{1}{\sqrt{y}}\int sinu^2 du."


By fundamental theorem of calculus ,every continuous real function f has an anti derivative but some ordinary looking continuous function have anti derivative that can not expressed in simple formula using more basic function. "I" looks like one of them.


"I" is called Fresnel integral which is a transcendental function.


"I" admits the following power series expansion that converges for all "x" .


"I=\frac{1}{\sqrt{y}}\int \sum_{n=0}^{n=\infin} \frac{ (-1)^n}{(2n+1)!}u^{2(2n+1)} du"



Where

"\text{sin}u=\sum_{n=0}^{n=\infin} \frac{(-1)^n}{(2n+1)!} u^{2n+1}"




"\implies I=\frac{1}{\sqrt{y}} \sum_{n=0}^{n=\infin} \frac{(-1)^n}{(2n+1)! (4n+3)}u^{4n+3}"




"\implies I=\frac{1}{\sqrt{y}} \sum_{n=0}^{n=\infin}\frac{(-1)^n}{(2n+1)!(4n+3)}x^{4n+3}y^{\frac{1}{2} (4n+3)}"